# MI-NeRF: Learning a Single Face NeRF from Multiple Identities

### v1

@@ B Additional Ablation Study
where $\bm{U}_{1},\bm{U}_{2},\bm{U}_{3}\in\mathbb{R}^{k\times d}$ and $\bm{C}\in\mathbb{R}^{o\times k}$. We set $o=k=d$. In this case, we capture third-order multiplicative interactions as well. The implicit representation $F_{\Theta}$ of the dynamic NeRF (see Eq. (3) of the main paper) would be:
$F_{\Theta}:(M(\bm{e},\bm{i},\bm{l}),\bm{x},\bm{v})\longrightarrow(\bm{c},\sigma)\;.$ (14)
@@
Figure 12: A visualization of our learned identity codes using t-SNE [36]. We notice that the identity codes capture meaningful information in terms of identity, e.g. different videos of Obama (bottom left) are clustered together, despite the different lighting.
C Additional Results
@@
Learned Identity Codes. Fig. 12 shows a visualization of our learned identity codes using t-SNE [36] on our final model with 100 identities. We notice that the identity codes capture meaningful information in terms of identity.
@@
Figure 13: Qualitative comparison of lip synced video synthesis by Wav2Lip [43], GeneFace [66], LipNeRF [8], and MI-NeRF. The original video is in English (1st column). The generated videos (columns 2-5) are lip synced to dubbed audio in Spanish.
@@
Additional Qualitative Results. Fig. 13 shows additional comparison with GeneFace [66] for lip synced video synthesis. GeneFace lacks in lip synchronization as also mentioned in Sec. 4.4. The GAN-based Wav2Lip [43] sometimes produces artifacts (e.g. see the mouth in the 2nd row). Our method achieves accurate lip syncing, while trained on multiple identities.
Fig. 15 shows additional results for our short-video personalization procedure, where MI-NeRF is adapted to an unseen identity (i.e. not seen as a part of the initial training set - see Sec. 3.4). Fig. 15(a) demonstrates qualitative results, when only a very short video of the new identity is available (1 or 3 seconds length). Please note that in this case we use a small number of consecutive frames, and thus a very small part of the expression space is covered. In contrast to NeRFace, MI-NeRF produces satisfactory lip shape and expression, as the multiplicative module has learned information from multiple identities during training. Correspondingly, Fig. 15(b) shows the LSE-D metric w.r.t different clip lengths of the target identity for the task of lip-syncing. Since AD-NeRF and LipNeRF are trained on a single identity, they learn audio-lip representations from only 3 or 20 seconds. On the other hand, MI-NeRF leverages information from multiple identities and leads to a more accurate lip synchronization.
(a) Short-Video Personalization
(b) LSE-D vs training clip length
Figure 15: Short-Video Personalization. (a) Qualitative comparison of MI-NeRF with the single-identity NeRFace [15], using a video of the target identity of only 1 or 3 seconds duration for adaptation. (b) LSE-D (lower the better) vs clip length of the target identity seen in training, computed on generated videos lip synced to dubbed audio.
@@
D Comparison with Faster Methods
We developed our multi-identity network and the proposed multiplicative modules based on the vanilla dynamic NeRF. We built upon dynamic NeRFs such as NeRFace [15], AD-NeRF [23] and LipNeRF [8]. However, we would like to note that recent methods propose techniques to reduce the training and rendering time of vanilla NeRFs. For example, INSTA [72] models the dynamic NeRF based on neural graphics primitives embedded around the parametric face model FLAME [31], requiring only 10 minutes training for an identity.
@@ D Comparison with Faster Methods
We specifically tried INSTA [72] and extended it to multiple identities. The original INSTA is a single-identity deformable NeRF, where an expression code $\bm{E}_{i}\in\mathbb{R}^{16}$ for an image $i$ conditions an MLP [72]. $\bm{E}_{i}$ is used for the input samples that correspond to the mouth region, and turned to $\bm{E}_{i}=\bm{1}$ otherwise. As in Eq. (1) of the main paper, we learned a multiplicative module $M(\bm{E}_{i},\bm{i})$ that conditions INSTA, with learnable identity codes $\bm{i}\in\mathbb{R}^{d}$ and $d=16$ in this case. We trained our multi-identity INSTA on our training set of 100 identities. We used their data pre-processing and FLAME fitting  [^4] https://github.com/Zielon/INSTA.
@@
Fig. 16 demonstrates the improvement when transferring a novel expression for INSTA trained on multiple identities, using our multiplicative module. However, we notice that INSTA produces some artifacts in the mouth interior (also mentioned by the authors [72]) that are not completely removed with our multi-identity optimization, and do not exist in our proposed MI-NeRF. With our multi-identity training, we similarly achieve a 90% decrease in the training time for 100 identities (see Sec. 4.3), compared to the single-identity INSTA. Specifically, INSTA requires about 10 minutes training for a single video, thus 1000 minutes total for 100 identities, and we train a multi-identity INSTA in less than 90 minutes for all 100 identities simultaneously.
@@
Figure 16: Extending other NeRF-based approaches to multiple identities. The single-identity INSTA [72] can be easily extended to a multi-identity INSTA by learning our proposed multiplicative module.
E Implementation Details
@@
In this section, we include additional implementation details. We closely follow the architecture and training details of NeRFace [15], AD-NeRF [23], and LipNeRF [8] that are all similar. An overview of the main hyper-parameters is given in Table 5. More specifically, our implementation is based on PyTorch [41]. We use Adam optimizer [27] with a learning rate that begins at $5\times 10^{-4}$ and decays exponentially to $5\times 10^{-5}$ during training. The rest of the Adam hyper-parameters are set at their default values ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$). For every gradient step, we march rays and sample points for a single randomly-chosen video frame. We march 2048 rays and sample 64 points per ray for the coarse volume and 128 points per ray for the fine volume (hierarchical sampling strategy [37]). We sample rays such that $95\%$ of them correspond to pixels inside the detected bounding box of the head [15]. We also assume that the last sample on each ray lies on the background and takes the corresponding RGB color [23, 8]. We use the original background per frame. Our MLP backbone consists of 8 linear layers with 256 hidden units each. The output of the backbone is fed to an additional linear layer to predict the density $\sigma$, and a 4-layer 128-unit wide branch to predict the RGB color $\bm{c}$ for every point. We use ReLU activations. Positional encodings are applied to both the input points $\bm{x}$ and the viewing directions $\bm{v}$, of 10 and 4 frequencies respectively. We do not apply smoothing on the expression parameters using a low pass filter, as proposed by [8]. Instead, we learn a self-attention, similarly to [23], that is applied to both the output of the multiplicative module and the latent codes after the first 200k iterations during training. This ensures smooth results in the final video synthesis, reducing any jitter. For 1-2 identities, we train our network for about 400k iterations (around 40 hours on a single GPU). For 10 or 20 identities, we need around 500k and 600k iterations correspondingly, and our generic MI-NeRF with 100 identities takes about 800k iterations (see Fig. 7 of the main paper). Further personalization requires another 50-80k iterations approximately for a target identity.
@@
To compute the visual quality metrics (PSNR, SSIM, LPIPS), we crop each frame around the face, using the face detector from 3DDFA [21, 20]. In this way, we evaluate the visual quality of the generated part only, ignoring the background that corresponds to the original one. To verify the speaker’s identity, we use the ACD metric, computing the cosine distance between the embeddings of the ground truth face and the generated face, extracted by InsightFace  [^5] https://github.com/deepinsight/insightface.
@@
F Limitations
An important factor is the 3DMM fitting (see Sec. 3.1 of the main paper) that is used as ground truth for the head pose and expression parameters. We use an existing optimization method [23]. However, this fitting can be noisy and the error can be propagated to the final generated videos. We address this by learning a self-attention, as mentioned in the implementation details above. Improving the face tracking further would be interesting to explore as future work. In addition, our network considers identity and expression, but does not learn interactions between other latent factors of variation, such as illumination or hair movement. Our model seems robust to different illuminations for the same identity (see Fig. 12). In the future, we plan to disentangle more latent factors of variation by extending our proposed high-degree interaction module $H$ to capture higher-degree interactions (see Sec. 3.2).
G Ethical Considerations
We would like to note the potential misuse of video synthesis methods. With the advances in neural rendering and generative models, it becomes easier to generate photorealistic fake videos of any identity. These can be used for malicious purposes, e.g. to generate misleading content and spread misinformation. Thus, it is important to develop accurate methods for fake content detection and forensics. Research on discriminative tasks has been investigated for several years by the community and there are certain guarantees and knowledge on how to build strong classifiers. However, this is only the first step towards mitigating the issue of fake content; further steps are required. A possible solution, that can be easily integrated in our work, is watermarking the generated videos [10], in order to indicate their origin. In addition, appropriate procedures must be followed to ensure fair and safe use of videos from a social and legal perspective.
H Dataset Details
In this section, we provide additional details for the dataset we used in our experiments. As mentioned in Sec. 4.1, we collected 140 talking face videos from publicly available datasets, which are commonly used in related works [23, 35, 8, 24, 19, 1, 14, 69, 63]. The detailed list of the 140 videos is as follows:
  • Standard videos used in related works, collected by [35] and [23]:
@@ H Dataset Details
  • PATS dataset [19, 1]:
@@
  • Actors from dataset proposed by LipNeRF [8]:
@@
  • Casual Conversations dataset [24]:
@@
Please note that most of these videos are from YouTube and the identities are public figures (e.g. politicians in HDTF dataset [69], famous actors in LipNeRF [8], comedians and professional YouTubers in PATS dataset [19, 1]  [^6] https://chahuja.com/pats/). The Casual Conversations dataset [24] includes talking videos of paid individuals who agreed to participate and opted-in for data use in ML. The participants are de-identified with unique numbers  [^7] https://ai.meta.com/datasets/casual-conversations-dataset/. The How2Sign dataset [14] is another dataset publicly available for research purposes that includes American Sign Language videos  [^8] https://how2sign.github.io/. TalkingHead-1kH [63] also includes videos from YouTube under permissive licenses only  [^9] https://github.com/tcwang0509/TalkingHead-1KH.
@@
For each video, we fit a 3DMM, as described in Sec. 3.1, in order to extract the corresponding pose and expression parameters of the identity per frame. We refer the interested reader to the work of [23]  [^10] https://github.com/YudongGuo/AD-NeRF
for more details in the data preprocessing.

### v2

@@ Appendix B Additional Ablation Study
where $\bm{U}_{1},\bm{U}_{2},\bm{U}_{3}\in\mathbb{R}^{k\times d}$ and $\bm{C}\in\mathbb{R}^{o\times k}$. We set $o=k=d$. In this case, we capture third-order multiplicative interactions as well. The implicit representation $F_{\Theta}$ of the dynamic NeRF (see Eq. 3 of the main paper) would be:
$F_{\Theta}:(M(\bm{e},\bm{i},\bm{l}),\bm{x},\bm{v})\longrightarrow(\bm{c},\sigma)$ (14)
@@
Figure 9: A visualization of our learned identity codes using t-SNE (Van der Maaten & Hinton, 2008). We notice that the identity codes capture meaningful information in terms of identity, e.g. different videos of Obama (bottom left) are clustered together, despite the different lighting.
Appendix C Additional Results
@@
Learned Identity Codes. Fig. 9 shows a visualization of our learned identity codes using t-SNE (Van der Maaten & Hinton, 2008) on our final model with 100 identities. We notice that the identity codes capture meaningful information in terms of identity.
@@
Figure 12: Qualitative comparison of lip synced video synthesis by Wav2Lip (Prajwal et al., 2020), GeneFace (Ye et al., 2023), LipNeRF (Chatziagapi et al., 2023), and MI-NeRF. The original video is in English (1st column). The generated videos (columns 2-5) are lip synced to dubbed audio in Spanish.
@@
Figure 14: Short-Video Personalization. Left: Qualitative comparison of MI-NeRF with the single-identity NeRFace (Gafni et al., 2020), using a video of the target identity of only 1 or 3 seconds duration for adaptation. Right: LSE-D (lower the better) vs clip length of the target identity seen in training, computed on generated videos lip synced to dubbed audio.
Additional Qualitative Results. Fig. 10 demonstrates additional qualitative results of facial expression transfer generated by MI-NeRF. Fig. 11 demonstrates how MI-NeRF can animate multiple identities under the same novel expression and novel view. Fig. 12 shows additional comparison with GeneFace (Ye et al., 2023) for lip synced video synthesis. GeneFace lacks in lip synchronization as also mentioned in Sec. 4.4. The GAN-based Wav2Lip (Prajwal et al., 2020) sometimes produces artifacts (e.g. see the mouth in the 2nd row). Our method achieves accurate lip syncing, while trained on multiple identities.
Fig. 14 shows additional results for our short-video personalization procedure, where MI-NeRF is adapted to an unseen identity (i.e. not seen as a part of the initial training set - see Sec. 3.4). Fig. 14 (left) demonstrates qualitative results, when only a very short video of the new identity is available (1 or 3 seconds length). Please note that in this case we use a small number of consecutive frames, and thus a very small part of the expression space is covered. In contrast to NeRFace, MI-NeRF produces satisfactory lip shape and expression, as the multiplicative module has learned information from multiple identities during training. Correspondingly, Fig. 14 (right) shows the LSE-D metric w.r.t different clip lengths of the target identity for the task of lip-syncing. Since AD-NeRF and LipNeRF are trained on a single identity, they learn audio-lip representations from only 3 or 20 seconds. On the other hand, MI-NeRF leverages information from multiple identities and leads to a more accurate lip synchronization.
@@
Appendix D Comparison with Faster Methods
We developed our multi-identity network and the proposed multiplicative modules based on the vanilla dynamic NeRF. We built upon dynamic NeRFs such as NeRFace (Gafni et al., 2020), AD-NeRF (Guo et al., 2021) and LipNeRF (Chatziagapi et al., 2023). However, we would like to note that recent methods propose techniques to reduce the training and rendering time of vanilla NeRFs. For example, INSTA (Zielonka et al., 2023) models the dynamic NeRF based on neural graphics primitives embedded around the parametric face model FLAME (Li et al., 2017), requiring only 10 minutes training for an identity.
@@ Appendix D Comparison with Faster Methods
We specifically tried INSTA (Zielonka et al., 2023) and extended it to multiple identities. The original INSTA is a single-identity deformable NeRF, where an expression code $\bm{E}_{i}\in\mathbb{R}^{16}$ for an image $i$ conditions an MLP (Zielonka et al., 2023). $\bm{E}_{i}$ is used for the input samples that correspond to the mouth region, and turned to $\bm{E}_{i}=\bm{1}$ otherwise. As in Eq. 1 of the main paper, we learned a multiplicative module $M(\bm{E}_{i},\bm{i})$ that conditions INSTA, with learnable identity codes $\bm{i}\in\mathbb{R}^{d}$ and $d=16$ in this case. We trained our multi-identity INSTA on our training set of 100 identities. We used their data pre-processing and FLAME fitting  [^4] https://github.com/Zielon/INSTA.
@@
Fig. 15 demonstrates the improvement when transferring a novel expression for INSTA trained on multiple identities, using our multiplicative module. However, we notice that INSTA produces some artifacts in the mouth interior (also mentioned by the authors (Zielonka et al., 2023)) that are not completely removed with our multi-identity optimization, and do not exist in our proposed MI-NeRF. With our multi-identity training, we similarly achieve a 90% decrease in the training time for 100 identities (see Sec. 4.3), compared to the single-identity INSTA. Specifically, INSTA requires about 10 minutes training for a single video, thus 1000 minutes total for 100 identities, and we train a multi-identity INSTA in less than 90 minutes for all 100 identities simultaneously.
@@
Figure 15: Extending other NeRF-based approaches to multiple identities. The single-identity INSTA (Zielonka et al., 2023) can be easily extended to a multi-identity INSTA by learning our proposed multiplicative module.
Appendix E Implementation Details
@@
In this section, we include additional implementation details. We closely follow the architecture and training details of NeRFace (Gafni et al., 2020), AD-NeRF (Guo et al., 2021), and LipNeRF (Chatziagapi et al., 2023) that are all similar. An overview of the main hyper-parameters is given in Table 5. More specifically, our implementation is based on PyTorch (Paszke et al., 2019). We use Adam optimizer (Kingma & Ba, 2014) with a learning rate that begins at $5\times 10^{-4}$ and decays exponentially to $5\times 10^{-5}$ during training. The rest of the Adam hyper-parameters are set at their default values ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$). For every gradient step, we march rays and sample points for a single randomly-chosen video frame. We march 2048 rays and sample 64 points per ray for the coarse volume and 128 points per ray for the fine volume (hierarchical sampling strategy (Mildenhall et al., 2020)). We sample rays such that $95\%$ of them correspond to pixels inside the detected bounding box of the head (Gafni et al., 2020). We also assume that the last sample on each ray lies on the background and takes the corresponding RGB color (Guo et al., 2021; Chatziagapi et al., 2023). We use the original background per frame. Our MLP backbone consists of 8 linear layers with 256 hidden units each. The output of the backbone is fed to an additional linear layer to predict the density $\sigma$, and a 4-layer 128-unit wide branch to predict the RGB color $\bm{c}$ for every point. We use ReLU activations. Positional encodings are applied to both the input points $\bm{x}$ and the viewing directions $\bm{v}$, of 10 and 4 frequencies respectively. We do not apply smoothing on the expression parameters using a low pass filter, as proposed by Chatziagapi et al. (2023). Instead, we learn a self-attention, similarly to Guo et al. (2021), that is applied to both the output of the multiplicative module and the latent codes after the first 200k iterations during training. This ensures smooth results in the final video synthesis, reducing any jitter. For 1-2 identities, we train our network for about 400k iterations (around 40 hours on a single GPU). For 10 or 20 identities, we need around 500k and 600k iterations correspondingly, and our generic MI-NeRF with 100 identities takes about 800k iterations (see Fig. 7 of the main paper). Further personalization requires another 50-80k iterations approximately for a target identity.
@@
To compute the visual quality metrics (PSNR, SSIM, LPIPS), we crop each frame around the face, using the face detector from 3DDFA (Guo et al., 2020; 2018a). In this way, we evaluate the visual quality of the generated part only, ignoring the background that corresponds to the original one. To verify the speaker’s identity, we use the ACD metric, computing the cosine distance between the embeddings of the ground truth face and the generated face, extracted by InsightFace  [^5] https://github.com/deepinsight/insightface.
@@
Appendix F Limitations
An important factor is the 3DMM fitting (see Sec. 3.1 of the main paper) that is used as ground truth for the head pose and expression parameters. We use an existing optimization method (Guo et al., 2021). However, this fitting can be noisy and the error can be propagated to the final generated videos. We address this by learning a self-attention, as mentioned in the implementation details above. Improving the face tracking further would be interesting to explore as future work. In addition, our network considers identity and expression, but does not learn interactions between other latent factors of variation, such as illumination or hair movement. Our model seems robust to different illuminations for the same identity (see Fig. 9). In the future, we plan to disentangle more latent factors of variation by extending our proposed high-degree interaction module $H$ to capture higher-degree interactions (see Sec. 3.2).
Appendix G Ethical Considerations
We would like to note the potential misuse of video synthesis methods. With the advances in neural rendering and generative models, it becomes easier to generate photorealistic fake videos of any identity. These can be used for malicious purposes, e.g. to generate misleading content and spread misinformation. Thus, it is important to develop accurate methods for fake content detection and forensics. Research on discriminative tasks has been investigated for several years by the community and there are certain guarantees and knowledge on how to build strong classifiers. However, this is only the first step towards mitigating the issue of fake content; further steps are required. A possible solution, that can be easily integrated in our work, is watermarking the generated videos (Chen et al., 2023), in order to indicate their origin. In addition, appropriate procedures must be followed to ensure fair and safe use of videos from a social and legal perspective.
Appendix H Dataset Details
In this section, we provide additional details for the dataset we used in our experiments. As mentioned in Sec. 4.1, we collected 140 talking face videos from publicly available datasets, which are commonly used in related works (Guo et al., 2021; Lu et al., 2021; Chatziagapi et al., 2023; Hazirbas et al., 2021; Ginosar et al., 2019; Ahuja et al., 2020; Duarte et al., 2021; Zhang et al., 2021; Wang et al., 2021b). The detailed list of the 140 videos is as follows:
  • Standard videos used in related works, collected by (Lu et al., 2021) and (Guo et al., 2021):
@@ Appendix H Dataset Details
  • PATS dataset (Ginosar et al., 2019; Ahuja et al., 2020):
@@
  • Actors from dataset proposed by LipNeRF (Chatziagapi et al., 2023):
@@
  • Casual Conversations dataset (Hazirbas et al., 2021):
@@
Please note that most of these videos are from YouTube and the identities are public figures (e.g. politicians in HDTF dataset (Zhang et al., 2021), famous actors in LipNeRF (Chatziagapi et al., 2023), comedians and professional YouTubers in PATS dataset (Ginosar et al., 2019; Ahuja et al., 2020)  [^6] https://chahuja.com/pats/). The Casual Conversations dataset (Hazirbas et al., 2021) includes talking videos of paid individuals who agreed to participate and opted-in for data use in ML. The participants are de-identified with unique numbers  [^7] https://ai.meta.com/datasets/casual-conversations-dataset/. The How2Sign dataset (Duarte et al., 2021) is another dataset publicly available for research purposes that includes American Sign Language videos  [^8] https://how2sign.github.io/. TalkingHead-1kH (Wang et al., 2021b) also includes videos from YouTube under permissive licenses only  [^9] https://github.com/tcwang0509/TalkingHead-1KH.
@@
For each video, we fit a 3DMM, as described in Sec. 3.1, in order to extract the corresponding pose and expression parameters of the identity per frame. We refer the interested reader to the work of (Guo et al., 2021)  [^10] https://github.com/YudongGuo/AD-NeRF for more details in the data preprocessing.